OpenStudy (anonymous):

can someone please help me with the rules for negative exponents 2^-2 (2)^-2 -2^2 (-2)^2 etc etc

OpenStudy (anonymous):

Something raised to a negative exponent is the same as it's reciprocal raised to a positive exponent.

OpenStudy (anonymous):

can u give me an example

OpenStudy (anonymous):

2^-2=1/2^2

OpenStudy (anonymous):

2^-2= (1/2) ^2 Those brackets are important if your numerator is not 1.

OpenStudy (anonymous):

What about in this situation? For the function y=x^2, compute the average rate of change between x=-3 and x=3 average rate of change is:

hero (hero):

Negative exponent really means to "inverse" the expression

hero (hero):

For example, the inverse of \[2^{-2}\] is \[\frac{1}{2^2}\] So it means to put 1 over the expression and remove the negative

hero (hero):

Average rate of change is just slope

OpenStudy (anonymous):

does that go for all numbers like 7^-13 would it be 1/7^13?

OpenStudy (anonymous):

yes

hero (hero):

yes, correct. The key to negative exponents is to realize that negative really means "inverse"

OpenStudy (anonymous):

i noticed that when you plug in numbers for example: What about in this situation? For the function y=x^2, compute the average rate of change between x=-3 and x=3 average rate of change is:

hero (hero):

As for finding the rate of change, you first have to find the y values Then create the points (x,y) and (x1,y1) Then use the slope formula

OpenStudy (anonymous):

it becomes a (-3)^2 rather than -3^2 why is that?

OpenStudy (anonymous):

because you need to include the negative in the square

hero (hero):

(-3)^2 means you are squaring -3 and the result will be positive -3^2 means you are squaring the 3 and the result will still be negative

hero (hero):

Once you find the y values, you will end up with points: (3,9) and (-3,9) When you calculate the slope (or the rate of change), you will realize that it is zero

OpenStudy (anonymous):

@Hero, I don't like how you used "inverse" to describe a reciprocal. Inverse means something completely different in relation to functions.

hero (hero):

replace inverse with reciprocal

OpenStudy (anonymous):

What about in this situation? For the function y=x^2, compute the average rate of change between x=-6 and x=-3 average rate of change is:-9 but i came up with POSITIVE 9 can you explain why ? I put (-6)^2 and (-3)^2 which would equal to positive 36 and 9 respectively

OpenStudy (anonymous):

(-6) - (-3) = -3

OpenStudy (anonymous):

formula is f(x2) - f(x1)/x2-x1

OpenStudy (anonymous):

so the -6 would not go first

OpenStudy (anonymous):

it doesn't matter which goes first, they're interchangeable.

hero (hero):

maybe I should have said Multiplicative Inverse

OpenStudy (anonymous):

when i plugged in -3-(-6) it came out positive 3

hero (hero):

Reciprocal = Multiplicative Inverse

OpenStudy (anonymous):

if you plugged in your -3 as your first x, you would have to plug in your 9 as your first y

OpenStudy (anonymous):

again, it doesn't matter which you plug in first, as long as the corresponding y is plugged in the same place.